OpenStudy (anonymous):

Find x^2 if x = whole root root 5 + 2 + whole root root 5 - 2/ whole root whole root 5 + 1.

OpenStudy (anonymous):

its 5

OpenStudy (anonymous):

Eh? Can you please explain with the steps.

OpenStudy (aum):

Can you post a screenshot of the question?

OpenStudy (anonymous):

How do I do that? I'm on my PC.

OpenStudy (anonymous):

4+1=5

OpenStudy (anonymous):

so its 5

OpenStudy (anonymous):

What exactly did you do, Kidz_Wavy? How did you come to the conclusion of '4 + 1 = 5'?

OpenStudy (anonymous):

lol 4 apples plus one apple=5

OpenStudy (aum):

If you press "Print Screen" (located near the top right corner on a Windows keyboard) it will save an image of the screen in memory. Then you can paste the image in a free program such as Irfanview and do some editing to remove extraneous info on your screen that you don't want people to see and then upload the image here.

OpenStudy (anonymous):

A curious child asked his mother: “Mommy, why are some of your hairs turning grey?” The mother tried to use this occasion to teach her child: “It is because of you, dear. Every bad action of yours will turn one of my hairs grey!” The child replied innocently: “Now I know why grandmother has only grey hairs on her head.” and that is 5...

OpenStudy (anonymous):

\[x = \sqrt{\sqrt{5} +2} + \sqrt{\sqrt{5} - 2} \div \sqrt{\sqrt{5 + 1}}\] find x^2

OpenStudy (anonymous):

its 5

OpenStudy (anonymous):

A curious child asked his mother: “Mommy, why are some of your hairs turning grey?” The mother tried to use this occasion to teach her child: “It is because of you, dear. Every bad action of yours will turn one of my hairs grey!” The child replied innocently: “Now I know why grandmother has only grey hairs on her head.”

OpenStudy (anonymous):

A curious child asked his mother: “Mommy, why are some of your hairs turning grey?” The mother tried to use this occasion to teach her child: “It is because of you, dear. Every bad action of yours will turn one of my hairs grey!” The child replied innocently: “Now I know why grandmother has only grey hairs on her head.”

OpenStudy (anonymous):

A curious child asked his mother: “Mommy, why are some of your hairs turning grey?” The mother tried to use this occasion to teach her child: “It is because of you, dear. Every bad action of yours will turn one of my hairs grey!” The child replied innocently: “Now I know why grandmother has only grey hairs on her head.”

OpenStudy (anonymous):

A curious child asked his mother: “Mommy, why are some of your hairs turning grey?” The mother tried to use this occasion to teach her child: “It is because of you, dear. Every bad action of yours will turn one of my hairs grey!” The child replied innocently: “Now I know why grandmother has only grey hairs on her head.”

OpenStudy (anonymous):

Yeah, I see how smart you are Kidz_Wavy.

OpenStudy (anonymous):

Is that better, aum?

OpenStudy (anonymous):

Can you answer me now? I certainly hope you can. I've an exam tomorrow. I don't think you really care but it's really important to me.

OpenStudy (anonymous):

k im sorry but the answer is 5

OpenStudy (aum):

\( \Large x = \frac{\sqrt{\sqrt{5} +2} + \sqrt{\sqrt{5} - 2}}{\sqrt{\sqrt{5 + 1}}} = \frac{\sqrt{\sqrt{5} +2} + \sqrt{\sqrt{5} - 2}}{\sqrt{\sqrt{6}}} \\ \Large x^2 = \frac{(\sqrt{5} +2) ~+~ (\sqrt{5} - 2) ~+~ 2\sqrt{\sqrt{5} +2} ~*~ \sqrt{\sqrt{5} -2}}{\sqrt{6}} = \frac{2\sqrt{5} ~+~ 2\sqrt{5 -4}}{\sqrt{6}} \\ \Large x^2 = \frac{2(\sqrt{5}+1)}{\sqrt{6}} = \frac{2\sqrt{6}(\sqrt{5}+1)}{6} = \frac{\sqrt{6}}{3}(\sqrt{5}+1) \)

OpenStudy (aum):

In the second line, the following algebraic identities were used: \(\Large (a+b)^2 = a^2 + b^2 + 2ab \\ \Large \sqrt{a}~*~\sqrt{b} = \sqrt{ab} \\ \Large (a+b)(a-b) = a^2 - b^2 \)